OpenStudy (anonymous):

plz help A new resturant opens in town. The first night 100 people came. For each successful night mfor a month 20more people showed up. How many poeple showed up on the 7th night?

OpenStudy (anonymous):

@dmezzullo @radar

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

i have to use the arithmetic series

OpenStudy (phi):

you could write down a few terms. 100, 120, 140,... in fact we could write down the number for all 7 nights, and get the answer that way. but if you have a series, there is a formula to find the 7th term do you know the formula?

OpenStudy (anonymous):

you could write down a few terms. 100, 120, 140,... in fact we could write down the number for all 7 nights, and get the answer that way. but if you have a series, there is a formula to find the 7th term do you know the formula?

OpenStudy (anonymous):

you really could

OpenStudy (anonymous):

the formula is sn[2a1+(n-1)d)]

OpenStudy (phi):

that looks peculiar how about this http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm

OpenStudy (phi):

you want the formula for the "nth term" not the sum of the terms

OpenStudy (anonymous):

yes thw nthg term

OpenStudy (anonymous):

the*

OpenStudy (phi):

and a formula has an = sign in it. Can you write down the formula for the nth term ?

OpenStudy (anonymous):

isnt it sn=n/2(a1+an)

OpenStudy (phi):

sn=n/2(a1+an) is the sum of the n terms. ( sn is short for SUM of n terms) it would give you 100+120+140+... but the question asks for How many poeple showed up on the 7th night they want just the number of people that showed up on that night. so you want the formula for the "nth term" (n will be 7 for this problem) see the left-hand box in http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm

OpenStudy (anonymous):

ok thanks

OpenStudy (phi):

The formula you want is \[ a_n = a_1 + (n-1)d \] d is the difference between 2 terms. Do you know what d is? a1 is the first term. They tell you it is 100 can you find a7 ?

OpenStudy (anonymous):

i got 220

OpenStudy (phi):

Let's see: d= 20 (they tell us this) a1=100 (they tell us this) n= 7 an= a1 + (n-1)*d replace the letters with numbers: a7= 100 + (7-1)*20 a7= 100+6*20 = 100+120= 220 yes, the 7th term is 220. as a check, we can do it the hard way: 100 <-- 1st day 120 140 160 180 200 220 <-- 7th day